An element X with an atomic mass of 60g/mol has density of 6.23g cm-3. If the edge length of its cubic unit cell is 400 pm, identify the type of cubic unit cell. Calculate the radius of an atom of this element.-?

Grade:12

2 Answers

Sunil Kumar FP
9 years ago
we have density d=ZM/a^3Na
Z=no of atom to calculate
M is the molecular mass
Na is the avagadro constant
a is the edge length
Z=(d*a^3Na)/M
=6.23*(4*10^-8)^3*6*10^23/60
=4
Therefore th unit cell is FCC.
and the length of the unit cell is sqrt2 a=4*radius=sqrt2/4*400pm=141.4pm
ankit singh
3 years ago
An element x with an atomic mass of 60 g/mol has density of 6.23 g/cm3. If the edge length of the unit cell is 400 pm, identify the type of cubic unit cell. ... ∴ Z=4 Hence, it is face-centred cubic unit cell.